# RDA-PSO: A computational method to quantify the diffusive dispersal of insects

### v1

@@ 3.6 Recapture of Diffusive Agents - Particle Swarm Optimization (RDA-PSO) model
The computational approach developed in this article provides a robust and accurate way of estimating the diffusion coefficient $D$ from temporal and spatial ratios associated with MRR data. The model is agent-based, treating mosquitoes as individual agents that perform a two-dimensional random walk in a region that mimics the conditions of the MRR experiment. The number of traps (we use the words trap or capture site interchangeably) in each zone around the release point is assumed to be known. The estimated value of $D$ is obtained via an optimization method that simultaneously fits the spatial (see Equation (5)) and temporal (see Equation (4)) recapture ratios. Specifically, the method aims to identify the parameters of the random walk that minimize the error $E$ between observed and simulated ratios, where $E$ is given in Equation (14) below. There are four such parameters:
@@
Capture sites are simulated as two concentric regions: the collection region is an inner disk of radius ${\tt q}\cdot{\tt h}$ in which walkers are captured, and the attracting region is an outer annulus of width ${\tt p}\cdot{\tt h}$, in which walkers are attracted toward the center of the inner disk. Whether or not a walker feels a nearby capture site is decided, for each step taken in the disk of radius ${\tt(p+q)}\cdot{\tt h}$ around the center of the trap, by a Bernoulli draw with probability of success ${\tt s}_{\tt e}$, equal to the efficiency of the associated trap (see Appendix B.2 for details).
@@
$\langle r^{2}\rangle_{1}=\text{number of steps traveled per unit of time}\cdot(\text{length of each step})^{2}=S\cdot({\tt k}\cdot{\tt h})^{2}.$
Setting $4D=S\cdot({\tt kh})^{2}$ leads to the above expression for $D$. Finding k is therefore equivalent to estimating $D$. We checked that changing $S$ in the RDA-PSO numerical method leads to changes in the optimal k that leave the estimated $D$ unchanged, thereby allowing us to set $S$ to the reasonable but arbitrary value of 12 steps per day (more details below).
@@
where $\{\sigma_{1},\dots,\sigma_{m}\}$ are the averaged simulated spatial ratios, $\{\tau_{1},\dots,\tau_{n}\}$ are the averaged simulated temporal ratios, $\{\sigma_{1}^{o},\dots,\sigma_{m}^{o}\}$ are the observed spatial ratios, and $\{\tau_{1}^{o},\dots,\tau_{n}^{o}\}$ are the observed temporal ratios. The inverse problem is solved in two steps: a grid search in the parameter space to find a suitable initial parameter range for k and q, followed by a particle swarm optimization (PSO) to find the parameter values that minimize $E$. The code allows the user to keep some parameters fixed (e.g. setting p = 0 and/or $s_{e}=100$) and perform the optimization on the remaining parameters.
@@ 3.6.1 Grid search
The purpose of the grid search is to obtain an overall description of the error landscape over a suitable region of the two-dimensional parameter space for k and q, with p = 0 and ${\tt s}_{\tt e}=100\%$. The forward problem is run at specific, uniformly distributed values of k and q, at which the error in Equation (14) is calculated. A smaller parameter range, in which the minimum error is expected to occur, is identified and serves as the initial k and q range for the PSO. The latter is performed with p ranging from 0 to the maximum allowable size so that traps do not overlap, and ${\tt s}_{\tt e}$ ranging from 0.1% to 100%. An example of error landscape associated with the two-parameter grid search is provided in Appendix D.
@@ 3.7 Implementation of the computational model
The RDA-PSO method is designed to be implemented on realistic settings representing actual MRR experiments. To make this possible, the forward problem is run at scale, with a value of h set at 0.1 meters. By default, mosquitoes are assumed to be released from a single point and recaptured in concentric zones around that point. The size of, and the number of capture sites contained in, each zone need to be provided. Capture sites may be traps (which typically only catch a small fraction of mosquitoes flying by) or houses, where mosquitoes are systematically collected with an aspirator. Appendix B provides more details on the image processing techniques used to digitize the study area, on the implementation of the random walk, including the locally biased motion in the attracting region of each site, and on the data collection.
@@ 4.1 Robustness of recapture ratios
The models discussed in this article use recapture ratios, defined in Section 3.1, as their input. We analyzed the robustness of these data via multiple runs of the forward model. As expected, recapture ratios vary with trap density. They are however independent from the number $N$ of mosquito released, as long as the latter is larger than a few hundred. They are robust to changes in trap efficiency if the trap density is low, which is the case for most real-world MRR experiments. For high trap density however, the recapture ratios depend on the value of se. Finally, up to a small uncertainty due to the randomness of each experiment, they appear to be independent of the exact location of the traps, as long as the number of traps in each collection zone is high enough and the trap efficiency remains low.
@@ 4.2 Properties of TC and ATC estimates
Appendix G reports on a systematic analysis of the performance of the time-corrected and area-and-time-corrected models, via a range of synthetic MRR experiments (for which we know the exact value of k). We consider situations where traps are evenly (leading to $D_{TC}=D_{ATC}$) or heterogeneously (where $D_{ATC}$ is relevant) distributed. When recapture ratios are robust against changes in trap location (while keeping the number of traps in each zone unchanged), the corresponding estimates of $D_{TC}$ and $D_{ATC}$ are also robust. In addition, these estimates become more accurate when temporal percentages are evaluated hourly (i.e. after each step of the random walk) instead of daily (which in the present case corresponds to every 12 steps). Consequently, the accuracy of these models is likely to be affected by the temporal resolution of MRR data. This is problematic since frequent sampling of traps or systematic insect aspiration inside collection houses is time-consuming. Moreover, as previously indicated, MRR experiments described in the literature report numbers of captured mosquitoes daily or less frequently. These shortcomings clearly limit the reliability of TC and ATC estimates. As documented below, the RDA-PSO method introduced in the present work is not hindered by such considerations.
@@ 4.3 Properties of RDA-PSO estimates
The RDA-PSO inverse method estimates the four parameter k, q, p, and ${\tt s}_{\tt e}$ from uncorrected temporal and spatial ratios (i.e. as defined in Equations (4) and (5) without adjustments). Uncertainty in the produced estimates results from arbitrariness of the initial conditions in parameter space for the zero-generation particles, from the stochastic nature of the forward model applied at each step of the method, as well as from the use of a finite number of generations. We confirmed that when applied to temporal and spatial ratios collected from virtual experiments, in which walkers spread with a known diffusion coefficient $D$, the RDA-PSO method recovers the actual value of $D$ with high accuracy. See Appendix D for an example.
To illustrate the applicability of the RDA-PSO method to real-life MRR data, we now consider two MRR experiments carried out in a small village on Hainan Island in China [76], and a third run in a suburban area of the city of Cairns in Australia [59].
4.3.1 Recapture ratios
In the Hainan MRR experiment [76], female mosquitoes were released at two different sites: one near the center and one at the periphery of the village. Details, including the village layout, are provided in Appendix E. The area covered by the Cairns MRR experiment is much larger (see Appendix F), with $r_{zone}=$  50 meters instead of 15 meters in Hainan. Reported recapture ratios for all three release sites are shown in Tables 1 (temporal ratios) and 2 (spatial ratios).
@@ 4.3.2 Parameter estimation
Figure 1 displays the results of the RDA-PSO method (with $N_{p}=36$ and $N_{g}=12$) applied to the Hainan and Cairns data shown in Tables 1 and 2. The search for the minimum was performed in the following regions of the parameter space.
(a) Hainan Center: ${\tt k}\in[15,55],\>{\tt q}\in[5,30],\>{\tt p}\in[0,15],\>s_{e}\in[0.001,1];$
(b) Hainan Edge: ${\tt k}\in[55,95],\>{\tt q}\in[5,40],\>{\tt p}\in[0,5],\>s_{e}\in[0.001,1];$
(c) Cairns: ${\tt k}\in[75,175],\>{\tt q}\in[5,40],\>{\tt p}\in[0,45],\>s_{e}\in[0.001,1].$
@@
Each panel of Figure 1 shows a scatter plot of the estimates of q, p, and ${\tt s}_{\tt e}$ as functions of the corresponding estimate of k. Also shown (purple dots) are the resulting values of the error (14). Outliers, associated with an error larger than the mean plus one half (for the Cairns data) or one standard deviation (for the Hainan data), were removed. In the first two panels, the returned values of p (squares) are near zero, which would justify the use of a simplified version of the RDA-PSO method in which p is set to zero, to better understand the role of the other parameters. This is not the case for the Cairns study. As illustrated in Appendices E and F, the general rule is that in the absence of p, the parameter q adjusts itself to produce a similar effect. Estimates of ${\tt s}_{\tt e}$ (multiplication signs) are relatively high, above 34.62%, 65.55%, and 32.12% in the first, second, and third panel respectively. Although some parameter estimates, especially for ${\tt s}_{\tt e}$, have large uncertainty, the range of returned values of k is much narrower. For the Hainan study, estimates of k are more dispersed for mosquitoes released at the edge of the village (note differences in the horizontal scale), which is to be expected since they first had to fly to the center of the village before being recaptured. In addition, the average value of k, and thus the estimation of $D_{RDA}$, is higher for mosquitoes released at the periphery (middle panel of Figure 1) than at the center (left panel of Figure 1) of the village, which also is to be expected.
@@
Figure 1: Results of the RDA-PSO method with tables 1 and 2 as input. Each returned optimal parameter 4-tuple, $({\tt k},{\tt q},{\tt p},{\tt s}_{\tt e})$, is represented as a set of 3 points, $({\tt k},{\tt q})$ (blue circles), $({\tt k},{\tt p})$ (red squares), and $({\tt k},{\tt s}_{\tt e})$ (yellow multiplication signs). The corresponding error is also plotted as a function of k (purple dots). Out of 500 parameter tuples obtained, those associated with an error one half standard deviation (Cairns) or one standard deviation (Hainan) above the mean are considered outliers and were removed. Left panel: Parameter estimation for the Hainan MRR experiment with release point at the center of the village. About 14.8% of the returned 4-tuples were considered outliers and are not plotted. Middle panel: Parameter estimation for the Hainan MRR experiment with release point at the periphery of the village (percentage of not-plotted outliers: 14.2%). Right panel: Parameter estimation for the Cairns MRR experiment (percentage of not-plotted outliers: 29.4%)
@@
We mentioned above that if p = 0, the variable q adjusts to compensate. Similarly, as se decreases, larger traps are needed in order to catch the same number of mosquitoes, and vice versa. In other words, the optimal values of the parameters describing the capture sites are interconnected and consequently not individually identifiable. Further details are provided in Appendices E and F (see Figures 8 and 10).
@@
4.4 Comparison of TC, ATC, MDT, and RDA-PSO estimates
This section compares the diffusion coefficients estimated with the time-corrected, area-and-time-corrected, MDT, and RDA-PSO models for the above MRR experiments.
Fitting the temporal and spatial capture ratios reported in each of Tables 1 and 2 into the time-corrected model (8) leads to estimates of the diffusion coefficient shown in the first column ($D_{TC}$) of Table 3. Since capture sites are not uniformly distributed across collection zones, we also show in column 2 ($D_{ATC}$) of Table 3 the area-and-time-corrected estimates of $D$. We note however that, contrary to the Cairns study, there is not sufficient information in the Hainan data [76] to correct the temporal ratios after having applied the correction factor to the spatial ratios. In other words, recapture numbers are reported per zone and per day in [59] but not in [76]. For comparison, the third column lists estimates of $D$ based on averages of reported values of the MDT: between 10 and 20 m for females released at the center of the village and close to 40 m for those released at the periphery in Hainan [76]; between 77.6 and 77.8 m in Cairns [59]. The last column shows estimates of $D$ obtained by applying the RDA-PSO method.
Table 3: Values of $D$ for the Hainan and Cairns MRR experiments. $D_{TC}$ and $D_{ATC}$ are calculated by fitting reported and corrected data, respectively, into the time-corrected and area-and-time-corrected models (8). Due to the limited information provided in [76], the temporal ratios used to estimate $D_{ATC}$ in the Hainan studies are the same as those used to estimate the corresponding $D_{TC}$. Estimates of $D_{MDT}$ are obtained from Equation (12) with temporal ratios given in Table 1 and MDT values equal to 15 m, 40 m, and 77.7 m respectively. The diffusion coefficient $D_{RDA}$ is obtained from the average of the k values shown in Figure 1, equal to 36.87, 71.99, 120.67 respectively.

### v2

@@ 3.5 Recapture of Diffusive Agents - Particle Swarm Optimization (RDA-PSO) model
The computational approach developed in this article provides a robust and accurate way of estimating the diffusion coefficient $D$ from temporal and spatial ratios associated with MRR data. The model is agent-based, treating mosquitoes as individual agents that perform a two-dimensional random walk in a region that mimics the conditions of the MRR experiment. Walkers do not naturally die (except when captured), but are allowed to leave the computational box, after which they do not return to the area of study. The number of traps (we use the words trap or capture site interchangeably) in each zone around the release point is assumed to be known. The estimated value of $D$ is obtained via an optimization method that simultaneously fits the spatial (see Equation (5)) and temporal (see Equation (4)) recapture ratios. Specifically, the method aims to identify the parameters of the random walk that minimize the error $E$ between observed and simulated ratios, where $E$ is given in Equation (14) below. There are four such parameters:
@@
The first parameter, k, is directly related to the diffusion coefficient, as shown below, while the last three parameters encapsulate properties of capture sites. For simplicity, the latter are simulated as two concentric regions: the collection region is an inner disk of radius ${\tt q}\cdot{\tt h}$ in which walkers are captured, and the attracting region is an outer annulus of width ${\tt p}\cdot{\tt h}$, in which walkers are attracted toward the center of the inner disk. Whether or not a walker feels a nearby capture site is decided, for each step taken in the disk of radius ${\tt(p+q)}\cdot{\tt h}$ around the center of the trap, by a Bernoulli draw with probability of success ${\tt s}_{\tt e}$, equal to the efficiency of the associated trap (see Appendix B.2 for details).
@@
[…]
Setting $4D=S\cdot({\tt kh})^{2}$ leads to the above expression for $D$. Finding k is therefore equivalent to estimating $D$. We checked that changing $S$ in the RDA-PSO numerical method leads to changes in the optimal k that leave the estimated $D$ unchanged, thereby allowing us to set $S$ to the reasonable but arbitrary value of 12 steps per day (see Section 3.6).
@@
where $\{\sigma_{1},\dots,\sigma_{m}\}$ are the averaged simulated spatial ratios, $\{\tau_{1},\dots,\tau_{n}\}$ are the averaged simulated temporal ratios, $\{\sigma_{1}^{o},\dots,\sigma_{m}^{o}\}$ are the observed spatial ratios, and $\{\tau_{1}^{o},\dots,\tau_{n}^{o}\}$ are the observed temporal ratios. The inverse problem is solved in two steps: a grid search in the parameter space to find a suitable initial parameter range for k and q, followed by a particle swarm optimization (PSO) to find the parameter values that minimize $E$. The code allows the user to keep some parameters fixed (e.g. setting p = 0 and/or $s_{e}=100$ %) and perform the optimization on the remaining parameters.
@@ 3.5.1 Grid search
The purpose of the grid search is to obtain an overall description of the error landscape over a suitable region of the two-dimensional parameter space for k and q, with p = 0 and ${\tt s}_{\tt e}=100\%$. The forward problem is run at specific, uniformly distributed values of k and q, at which the error in Equation (14) is calculated. An example of error landscape associated with synthetic data is provided in Appendix D. This error landscape clearly identifies a global minimum (see Figure 5), which corresponds to the known k and q parameters. Similarly, in the case of field data, a global minimum was always observed in the examples we considered, as shown in Appendices E and F.
However, to mimic real-world uncertainties and include the effect of the other two parameters when analyzing field experiments, we do not rely on this straightforward search method and instead develop a better optimization technique, as shown below. Specifically, the grid search serves to identify a smaller parameter range, in which the global minimum error is expected to occur, and provides the initial k and q range for the PSO. The latter is then performed with p ranging from 0 to the maximum allowable size so that traps do not overlap, and ${\tt s}_{\tt e}$ ranging from 0.1% to 100%.
@@ 3.6 Implementation of the computational model
The RDA-PSO method is designed to be implemented on realistic settings representing actual MRR experiments. To make this possible, the forward problem is run at scale, with a value of h set at 0.1 meter. By default, mosquitoes are assumed to be released from a single point and recaptured in concentric zones around that point. The size of, and the number of capture sites contained in, each zone need to be provided. Capture sites may be traps (which typically only catch a small fraction of mosquitoes flying by) or houses, where mosquitoes are systematically collected with an aspirator. Appendix B provides more details on the image processing techniques used to digitize the study area, on the implementation of the random walk, including the locally biased motion in the attracting region of each site, and on the data collection.
@@
Table 1 below summarizes the computational setup and the data needed for the PSO implementation.
Table 1: RDA-PSO parameters. The default parameters may be adjusted by the user. The experimental constants/data depend on the field experiment that is being analyzed. The PSO-configuration parameters are set as a compromise between computational cost and result accuracy. See Appendix C for further information on the PSO configuration.
parameter label/value
Default parameters mesh size of the physical landscape h = 0.1 (in m)
number of steps agents take per unit time S = 12
Experimental constants/data number of mosquitoes released NN
number of collection days $n$
width of each concentric zone $r_{zone}$ (from maps)
release and capture site locations (from maps)
recapture ratios $\tau_{i}$ and $\sigma_{j}$
PSO-configuration parameters number of swarming particles $N_{p}=36$
number of generations $N_{g}=12$
[…]
@@ 4.1 Robustness of recapture ratios
The models discussed in this article use recapture ratios, defined in Section 3.1.2, as their input. We analyzed the robustness of these data via multiple runs of the forward model. As expected, recapture ratios vary with trap density. They are however independent from the number $N$ of mosquito released, as long as the latter is larger than a few hundred. They are robust to changes in trap efficiency if the trap density is low, which is the case for most real-world MRR experiments. For high trap density however, the recapture ratios depend on the value of se. For example, if traps are uniformly distributed over the region $\mathcal{R}$ and there is one trap in the first zone (of radius $15$ m), the resulting recapture ratios are similar regardless of trap efficiency. However, if the density of traps is multiplied by 4, recapture ratios vary significantly with the chosen trap efficiency se. Details are provided in Appendix H, where we remark that the latter situation would correspond to a badly designed MRR experiment, since many of the traps located in the outer zones would not catch any mosquitoes at all, and the initial pool of mosquitoes would be essentially exhausted after a few days. Finally, up to a small uncertainty due to the randomness of each experiment, recapture ratios appear to be independent of the exact location of the traps, as long as the number of traps in each collection zone is high enough and the trap efficiency remains low.
@@ 4.2 Properties of TC and ATC estimates
Appendix G reports on a systematic analysis of the performance of the time-corrected and area-and-time-corrected models, via a range of synthetic MRR experiments (for which we know the exact value of k). We consider situations where traps are evenly (leading to $D_{TC}=D_{ATC}$) or heterogeneously (where $D_{ATC}$ is relevant) distributed. When recapture ratios are robust against changes in trap location (while keeping the number of traps in each zone unchanged), the corresponding estimates of $D_{TC}$ and $D_{ATC}$ are also robust. In addition, these estimates become more accurate when temporal percentages are evaluated hourly (i.e. after each step of the random walk) instead of daily (which in the present case corresponds to every 12 steps). Some of these results are summarized in Table 2, for uniform trap distributions. Results for non-uniformly distributed traps are presented in Appendix G.2, where it is shown that, on average, estimates of $D_{ATC}$, based on corrected temporal and spatial ratios and hourly data collection, tend to underestimate the known value $D$.
Table 2: Diffusion coefficient with different temporal resolutions of MRR data. Traps are uniformly distributed over the region $\mathcal{R}$, so $D_{TC}=D_{ATC}$. Low trap density corresponds to 1 trap in the first zone, while high trap density corresponds to 4 traps. Trap efficiency is set at 3%. More setup information is provided in Appendix G, and values listed are part of the data used to obtain figure 11.
exact $D$ $D_{ATC}(=D_{TC})$ [min, max]
[…]
low trap density daily capture 36.75 33 [22.21, 41.3]
hourly capture 36.29 [24.65, 42.82]
high trap density daily capture 30.79 [26.93, 33.75]
hourly capture 35.59 [32.31, 38.42]
[…]
Regardless of trap distribution, estimates of $D$ are affected by the temporal resolution (daily vs. hourly) of MRR data. This is problematic since frequent sampling of traps or systematic insect aspiration inside collection houses is time-consuming. Moreover, as previously indicated, MRR experiments described in the literature report numbers of captured mosquitoes daily or less frequently. These shortcomings clearly limit the reliability of TC and ATC estimates. As documented below, the RDA-PSO method introduced in the present work is not hindered by such considerations.
@@ 4.3 Properties of RDA-PSO estimates
The RDA-PSO inverse method estimates the four parameter k, q, p, and ${\tt s}_{\tt e}$ from uncorrected temporal and spatial ratios (i.e. as defined in Equations (4) and (5) without adjustments). Estimates of the diffusion coefficient $D$ depend directly on estimates of k, as shown in equation (13). Uncertainty in the produced estimates results from arbitrariness of the initial conditions in parameter space for the zero-generation particles, from the stochastic nature of the forward model applied at each step of the method, as well as from the use of a finite number of generations. We confirmed that when applied to temporal and spatial ratios collected from virtual experiments, in which walkers spread with a known diffusion coefficient $D$, the RDA-PSO method recovers the actual value of k and thus of $D$ with high accuracy. See Appendices D and G for two examples. The remaining three parameters, q, p, and ${\tt s}_{\tt e}$, are all features of the capture sites. Our numerical explorations indicate that these parameters are not individually identifiable and that their RDA-PSO estimates adjust themselves relative to one another in predictable ways. Importantly, we note that each estimation of the triplet (q, p, and ${\tt s}_{\tt e}$) produced by the RDA-PSO method is associated with a reliable estimate of the diffusion coefficient, which is therefore identifiable.
4.3.1 Field experiments
To illustrate the applicability of the RDA-PSO method to real-life MRR data, we now consider two MRR experiments carried out in a small village on Hainan Island in China [80] (we only consider the female cohorts from this study), and a third run in a suburban area of the city of Cairns in Australia [63].
In the Hainan MRR experiment [80], female mosquitoes were released at two different sites: one near the center and one at the periphery of the village. Details, including the village layout, are provided in Appendix E. The area covered by the Cairns MRR experiment is much larger (see Appendix F), with $r_{zone}=$  50 meters instead of 15 meters in Hainan. Reported recapture ratios for all three release sites are shown in Tables 3 (temporal ratios) and 4 (spatial ratios).
@@ 4.3.2 Parameter estimation
Figure 1 displays the results of the RDA-PSO method (with $N_{p}=36$ and $N_{g}=12$) applied to the Hainan and Cairns data shown in Tables 3 and 4. The search for the minimum was performed in the following regions of the parameter space.
$\text{(a) Hainan Center: }{\tt k}\in[15,55],\>{\tt q}\in[5,30],\>{\tt p}\in[0,15],\>s_{e}\in[0.001,1];$
$\text{(b) Hainan Edge: }{\tt k}\in[55,95],\>{\tt q}\in[5,40],\>{\tt p}\in[0,5],\>s_{e}\in[0.001,1];$
$\text{(c) Cairns: }{\tt k}\in[75,175],\>{\tt q}\in[5,40],\>{\tt p}\in[0,45],\>s_{e}\in[0.001,1].$
@@
Each panel of Figure 1 shows a scatter plot of the estimates of q, p, and ${\tt s}_{\tt e}$ as functions of the corresponding estimate of k. Also shown (purple dots) are the resulting values of the error (14). Outliers, associated with an error larger than the mean plus one half (for the Cairns data) or one standard deviation (for the Hainan data), were removed. In the first two panels, the returned values of p (squares) are near zero, which would justify the use of a simplified version of the RDA-PSO method in which p is set to zero, to better understand the role of the other parameters. This is not the case for the Cairns study. As illustrated in Appendices E and F, the general rule is that in the absence of p, the parameter q adjusts itself to produce similar capture ratios. Estimates of ${\tt s}_{\tt e}$ (multiplication signs) are relatively high, above 34.62%, 65.55%, and 32.12% in the first, second, and third panel respectively. Although some parameter estimates, especially for ${\tt s}_{\tt e}$, have large uncertainty, the range of returned values of k is much narrower. For the Hainan study, estimates of k are more dispersed for mosquitoes released at the edge of the village (note differences in the horizontal scale), which is to be expected since they first had to fly to the center of the village before being recaptured. In addition, the average value of k, and thus the estimation of $D_{RDA}$, is higher for mosquitoes released at the periphery (middle panel of Figure 1) than at the center (left panel of Figure 1) of the village, which also is to be expected.
@@
Figure 1: Results of the RDA-PSO method with tables 3 and 4 as input. Each returned optimal parameter 4-tuple, $({\tt k},{\tt q},{\tt p},{\tt s}_{\tt e})$, is represented as a set of 3 points, $({\tt k},{\tt q})$ (blue circles), $({\tt k},{\tt p})$ (red squares), and $({\tt k},{\tt s}_{\tt e})$ (yellow multiplication signs). The corresponding error is also plotted as a function of k (purple dots). Out of 500 parameter tuples obtained, those associated with an error one half standard deviation (Cairns) or one standard deviation (Hainan) above the mean are considered outliers and were removed. Left panel: Parameter estimation for the Hainan MRR experiment with release point at the center of the village. About 14.8% of the returned 4-tuples were considered outliers and are not plotted. Middle panel: Parameter estimation for the Hainan MRR experiment with release point at the periphery of the village (percentage of not-plotted outliers: 14.2%). Right panel: Parameter estimation for the Cairns MRR experiment (percentage of not-plotted outliers: 29.4%)
@@
We mentioned above that if p = 0, the variable q adjusts to compensate. Similarly, as se decreases, larger traps are needed in order to catch the same number of mosquitoes, and vice versa. In other words, the optimal values of the parameters describing the capture sites are interconnected and consequently not individually identifiable. However, the combination of optimal parameters, q, p, se, is always associated with an optimal k parameter that falls within the normal distribution mentioned above, leading to a reliable estimate for $D_{RDA}$. Further details are provided in Appendices E and F (see Figures 8 and 10 for more information on the interconnectedness of trap-related parameters).
@@
4.4 Comparison of MDT, TC, ATC, and RDA-PSO estimates
This section compares the diffusion coefficients estimated with the MDT, time-corrected, area-and-time-corrected, and RDA-PSO models for the above MRR experiments. All estimates are summarized in Table 5.
Table 5: Values of $D$ for the Hainan and Cairns MRR experiments. Estimates of $D_{MDT}$ are obtained from Equation (7) with temporal ratios given in Table 3 and MDT values equal to 15 m, 40 m, and 77.7 m respectively. $D_{TC}$ and $D_{ATC}$ are calculated by fitting reported and corrected data, respectively, into the time-corrected and area-and-time-corrected models (10). Due to the limited information provided in [80], the temporal ratios used to estimate $D_{ATC}$ in the Hainan studies are the same as those used to estimate the corresponding $D_{TC}$. The diffusion coefficient $D_{RDA}$ is obtained from the average of the k values shown in Figure 1, equal to 36.87, 71.99, 120.67 respectively.